A rectangular gutter is formed by bending a 24 cm wide sheet of metal as shown. As a whole number, what is the height of a the gutter of maximum capacity?


1
Expert's answer
2022-06-06T08:32:20-0400

Let "x="the height of the gutter. Then the width of the gutter will be "24-2x."

The capacity is

"S=S(x)=x(24-2x), 0<x<12"

The quadratic function has the absolute maximum at


"x=-\dfrac{24}{2(-2)}=6"

The height of a the gutter of maximum capacity is 6 cm.